# Second multicols Demo A Study of thin relativistic viscose accretion disk around a distorted kerr black hole (DKB)

### v1

@@ Abstract
In this paper, we analyze the distorted Kerr black hole (DKB) within the framework of general relativity using an axisymmetric solution of the Einstein equations. We consider the Kerr black hole in an external gravitational field up to the quadrupole moment and discuss the key aspects of black hole accretion disk theory.
Our findings indicate that the presence of a quadrupole moment significantly influences the radiation emitted from the accretion disk. While the location of the innermost stable circular orbit (ISCO) remains largely unchanged, the magnitude of the radiation flux, as well as the shape, orientation, and energy distribution of the accretion disk, are affected. The direction of distortion of the event horizon determines whether the disk becomes more oblate or prolate, impacting observed variations in maximum height, position, and temperature. Furthermore, the quadrupole moment alters the geometry of the black hole’s spacetime, which can influence the efficiency of energy extraction from the black hole’s spin—an important factor in powering emissions from accretion disks. Additionally, we investigated the effects of varying the viscosity coefficient on the behavior of the DKB. We also examined how rotation influences the dynamics of the DKB.
@@ 1 Introduction
A huge part of the light we receive from the depths of the universe reaches us from systems in which accretion played a major role. The maximum luminosity of the disk is obtained from the Eddington luminosity relation, where erg/s. Most analyzes of accretion disk models assume a steady and axisymmetric mode of matter accretion into a black hole. In these models, all physical quantities depend on only two spatial coordinates: the radial distance from the center r and the vertical distance from the plane of equatorial symmetry z. Most of the studied models assume that the disk is not vertically thick and the thin disk. In thin disks, inside the matter distribution [1]. In models of thin accretion disks Eddington’s critical mass , the first studies were done by Shakura and Sunyaev [2]. The black hole distorts the space around it, which warps images of stars lined up almost directed behind it.
@@
Relativistic disks were first described by Bardeen, Press and Tokolsky in 1972 that outlines a technique for analyzing physical processes around rotating black holes [3] and Novika and Tourne in 1973 [4]. For the first time in 1965, Doroshkevich considered a Schwarzschild black hole in a quadrupole gravitational field with a certain horizon [5]. Geroch and Hartl then did further analytical work on the warped black hole in 1982. They are obtained all exact solutions of Einstein’s equation that represent static, axisymmetric black holes distorted by an external matter distribution [6]. In 2001, Fairhurst, Krishnan and Marolf presented new solutions to the Einstein-Maxwell equations representing a class of charged distorted black holes. These solutions are static-axisymmetric and generalizations of the distorted black hole solutions [7]. They discussed thermodynamics of such distorted black holes like Stoytcho S Yazadjiev [8]. Shoom studied the interior of distorted stationary rotating black holes on the example of a Kerr black hole distorted by external static and axisymmetric mass distribution [9] and studied geodesic motion around a distorted Schwarzschild black hole [10]. Shohreh Abdolrahimi has investigated the properties of the ergo region and the location of the curvature singularities for the Kerr black hole distorted by the gravitational field of external sources. The particular cases of quadrupole and octupole distortion are studied in detail [11]. Grover et al studied the local shadow of the Schwarzschild black hole with a quadrupole distortion and the influence of the external gravitational field on the photon dynamics [12]. A quadrupole refers to a specific distribution pattern in space. It describes a distribution of mass, energy, or other physical quantities where the center of mass is balanced but the distribution is not perfectly uniform. In essence, quadrupole moments are a mathematical concept used to describe non-uniform distributions in space. The quadrupole moment significantly impacts the geometry of the black hole, leading to a departure from spherical symmetry and affecting the paths of particles and light around the black hole. In the simulations and analytical models of accretion disks, the energy content is considered in such a way that its effect on the space-time geometry is negligible.
@@
This paper focuses on the distorted Kerr black hole (DKB) within the framework of general relativity, employing an axisymmetric solution of the Einstein equations. We first consider the Kerr black hole in an external gravitational field up to the quadrupole, as constructed by Abdolrahimi [11]. Novikov and Thorne [4] provided a detailed analysis of the Kerr spacetime, while Abramowicz [1] delved into the key aspects of black hole accretion disk theory. These works highlight how accretion disks can reveal unique signatures of strong gravity, including the event horizon and the innermost stable circular orbit (ISCO).
In the third section of this paper, we delve into the metric assumptions and constants. We leverage the work of Faraji [13], who constructed a standard relativistic thin disk around a distorted Schwarzschild black hole up to the quadrupole. We then proceed to solve the conservation equations in the DKB metric. As the mass of the black hole is comparable to that of the sun (), the effect of self-gravity is negligible.
we initially focused on the variations in the quadrupole moment while keeping the viscosity coefficient constant for both rotating and non-rotating configurations. Following this, we investigated a range of viscosity coefficient values for both distorted and undistorted states, ensuring that the rotational parameter remained fixed. Next, we assigned specific viscosity coefficient values to the undistorted and distorted Kerr black holes, making necessary adjustments to the rotational parameter. Through detailed graphical representations, we performed a comprehensive analysis of their behaviors and the implications of these modifications. Our findings contribute to a deeper understanding of the dynamics involved in black hole configurations and their physical characteristics.
The results of our analysis, focusing on the thermodynamic effects of mass distribution around this scenario, are presented in the form of graphs and detailed explanations. We chose to study the DKB due to the significance of rotation in black holes and the accuracy of the model in representing the motion of the accretion disk and the central black hole, which aligns with real-world observations.
Further details and explanations of the calculations are provided in Appendix .
2 Thin Disk
Distortions are found in binary systems where the black hole forms an accretion disk with its compantion. They can be investigated numerically and analytically. many works have been done to numerically investigate such systems and merge black holes [20].
Also, some black holes can be modeled using axisymmetric and stationary solutions. In these cases, precise solutions for distorted black holes are examined.
The Kerr metric describes the geometry of spacetime around a charge-free rotating axisymmetric black hole, which is an exact solution of Einstein’s field equations of general relativity. The DKB metric is given in the following form [11], [21]:
@@ 2 Thin Disk
We denote by the Legendre polynomials, which depends on the argument xy/R in all the expressions. The solution is represented in the prolate spheroidal coordinates, m>0, and =q , are real constants. The horizon of the DKB is located at x = 1. So:
@@
The following assumptions are considered for the calculations and conditions related to the DKB and the accretion disk:
- A standard relativistic thin disk around a Kerr black hole is assumed [22] .
- The disk is geometrically thin in which H is half thickness of the disk.
- It’s cold which means
- Optically thick which means that the mean free path of photons is short.
- We focused on the region beyond the horizon that exhibits both Axial and mass symmetry, It is rotating , and this distortion maybe related to the outer parts of the accretion disk.
@@
- Innermost Stable Circular Orbit (ISCO): This is radius within which free circular orbital motion is not possible. We did not consider any torque for the inner edge of the disk [4].
In [12] it is shown that the ISCO exists only for
@@ 2.1 Conservation Equations
The equations are:
1. Describes the behavior of fluid flow in a radial direction under relativistic conditions:
@@
2. Particles conservation equation:
@@
3. The energy conservation equation is given as:
@@
3 Summary and Results
We worked on the metric equations of curved space time and simplified them, ultimately discovering new algebraic equations as a result. The obtained results are depicted here, where we compare the plots of Kerr’s distorted and undistorted spacetime. Since the distorted solutions are only applicable in a limited region near the horizon (we consider the horizon to be at ), we focus solely on the inner portion of the ISCO disc within this spacetime. The accuracy of our exact solution depends on the chosen quadrupole. Additionally, we assume that electron scattering and radiation pressure are the prevailing factors. The constants involved are as follows: (All constants are in the CGS system)
@@ 3 Summary and Results
Now, from [4], we have written the equations for DKB:
In the following sections, based on equations 53 to 58, we present graphs illustrating the flux, surface density, half-thickness, temperature, pressure, and viscosity as functions of the radial coordinates (distance from the black hole). These graphs account for various states specified in the captions of the figures:
![[Uncaptioned image]](x1.png)
![[Uncaptioned image]](x2.png)
@@
![[Uncaptioned image]](x3.png)
![[Uncaptioned image]](x4.png)
In figures 3 and 4, the radial coordinate x is plotted against the surface density . At x = 3, we observe the maximum surface density without rotation at the initial point. However when rotation is applied, this initial maximum shifts closer to the horizon. When we compare each quadrupole moment q in both states, we notice that in the compare rotation state (Kerr), the growth of surface density occurs more rapidly. This indicates the surface density decreases with the distance from the horizon, which we expect, but by adding rotation It is increasing that needs to be investigated more so that in higher positions it should tend to zero in order to get the correct result.
@@
![[Uncaptioned image]](x13.png)
![[Uncaptioned image]](x14.png)
![[Uncaptioned image]](x15.png)
@@
![[Uncaptioned image]](x16.png)
![[Uncaptioned image]](x17.png)
![[Uncaptioned image]](x18.png)
@@
![[Uncaptioned image]](x19.png)
![[Uncaptioned image]](x20.png)
![[Uncaptioned image]](x21.png)
@@
![[Uncaptioned image]](x22.png)
![[Uncaptioned image]](x23.png)
![[Uncaptioned image]](x24.png)
@@
![[Uncaptioned image]](x25.png)
![[Uncaptioned image]](x26.png)
![[Uncaptioned image]](x27.png)
@@
![[Uncaptioned image]](x28.png)
![[Uncaptioned image]](x29.png)
![[Uncaptioned image]](x30.png)
@@
![[Uncaptioned image]](x31.png)
![[Uncaptioned image]](x32.png)
![[Uncaptioned image]](x33.png)
@@
![[Uncaptioned image]](x34.png)
![[Uncaptioned image]](x35.png)
![[Uncaptioned image]](x36.png)
@@
![[Uncaptioned image]](x37.png)
![[Uncaptioned image]](x38.png)
![[Uncaptioned image]](x39.png)
@@
![[Uncaptioned image]](x40.png)
![[Uncaptioned image]](x41.png)
![[Uncaptioned image]](x42.png)
@@
![[Uncaptioned image]](x43.png)
![[Uncaptioned image]](x44.png)
![[Uncaptioned image]](x45.png)
@@
![[Uncaptioned image]](x46.png)
![[Uncaptioned image]](x47.png)
![[Uncaptioned image]](x48.png)
@@ 4 Discussion and conclusion
In this article, we first examined the changes in the quadrupole moment at a fixed alpha for both rotating and non-rotating states. Next, we varied alpha for quadrupoles with values of zero, positive, and negative, while keeping the rotation parameter at a positive value less than zero. Subsequently, we fixed alpha for the zero, positive, and negative quadrupoles and adjusted the rotation parameter. By plotting all relevant graphs, we analyzed their behaviors and effects.
@@
b) -x graph: An increase in rotation (a) and a decrease in viscosity () results in an additional peak in the surface density-radial coordinate (-x) graph. A positive quadrupole accelerates the growth of the graph, while a negative quadrupole exhibits slower growth.
@@
d) T-x graph: Increasing rotation (a) and viscosity () elevates the average temperature in the temperature-radial coordinate (T-x) graph. Changes in quadrupole primarily affect the slope of the graph over time, with no noticeable impact at the initial stages.

### v3

@@ Abstract
In this paper, we analyze a thin disk around the distorted Kerr black hole (DKB) within the framework of general relativity using an axisymmetric solution of the Einstein equations. We consider this accretion disk around the Kerr black hole in an external gravitational field up to the quadrupole moment and discuss the key aspects of black hole accretion disk theory.
Our findings indicate that the presence of a quadrupole moment significantly influences the radiation emitted from the accretion disk. While the location of the innermost stable circular orbit (ISCO) remains largely unchanged, the magnitude of the radiation flux, as well as the shape, orientation, and energy distribution of the accretion disk, are affected. The direction of distortion of the event horizon determines whether the disk becomes more oblate or prolate, impacting observed variations in maximum height, position, and temperature. Furthermore, the quadrupole moment alters the geometry of the black hole’s spacetime, which can influence the efficiency of energy extraction from the black hole’s spin—an important factor in powering emissions from accretion disks. We obtain dynamical quantities around a distorted rotating black hole disk. Additionally, we examine how rotation influences the dynamics of the DKB. We also investigate the effects of varying the viscosity coefficient on the behavior of the DKB.
@@ 1 Introduction
A huge part of the light we receive from the depths of the universe reaches us from systems in which accretion played a major role. The luminosity of the disk is obtained from the Eddington luminosity relation, where erg/s. Most analyzes of accretion disk models assume a steady and axisymmetric mode of matter accretion into a black hole. In these models, all physical quantities depend on only two spatial coordinates: the radial distance from the center r and the vertical distance from the plane of equatorial symmetry z. Most of the studied models assume that the disk is not vertically thick and the thin disk. In thin disks, inside the matter distribution [1]. In models of thin accretion disks where =16/ and Eddington’s critical mass , the first studies were done by Shakura and Sunyaev [2]. The black hole distorts the space around it, which warps images of stars lined up almost directed behind it.
@@
Bardeen, Press and Teukolsky in 1972 developed techniques for studying perturbations and relativistic processes around rotating black holes [3]. The first formal model of relativistic thin accretion disks was introduced by Novika and Torne in 1973 [4]. For the first time in 1965, Doroshkevich considered a Schwarzschild black hole in a quadrupole gravitational field with a certain horizon [5]. Geroch and Hartl then did further analytical work on the warped black hole in 1982. They are obtained all exact solutions of Einstein’s equation that represent static, axisymmetric black holes distorted by an external matter distribution [6]. In 2001, Fairhurst, Krishnan and Marolf presented new solutions to the Einstein-Maxwell equations representing a class of charged distorted black holes. These solutions are static-axisymmetric and generalizations of the distorted black hole solutions [7]. They discussed thermodynamics of such distorted black holes like Stoytcho S Yazadjiev [8]. Shoom studied the interior of distorted stationary rotating black holes on the example of a Kerr black hole distorted by external static and axisymmetric mass distribution [9] and studied geodesic motion around a distorted Schwarzschild black hole [10]. Shohreh Abdolrahimi has investigated the properties of the ergo region and the location of the curvature singularities for the Kerr black hole distorted by the gravitational field of external sources. The particular cases of quadrupole and octupole distortion are studied in detail [11]. Grover et al studied the local shadow of the Schwarzschild black hole with a quadrupole distortion and the influence of the external gravitational field on the photon dynamics [12]. A quadrupole specifically refers to the second-order term in a multipole expansion, characterizing deviations from spherical symmetry in the mass-energy distribution. While the centre of mass is balanced and it measures asymmetries in the gravitational field. In essence, quadrupole moments are a mathematical concept used to describe non-uniform distributions in space. The quadrupole moment significantly impacts the geometry of the black hole, leading to a departure from spherical symmetry and affecting the paths of particles and light around the black hole. In the simulations and analytical models of accretion disks, the energy content is considered in such a way that its effect on the space-time geometry is negligible.
@@
This paper focuses on the DKB within the framework of general relativity, employing an axisymmetric solution of the Einstein equations. In section 2, we consider the Kerr black hole in an external gravitational field up to the quadrupole, as constructed by Abdolrahimi [11]. Novikov and Thorne [4] provided a detailed analysis of the Kerr spacetime, while Abramowicz [1] delved into the key aspects of black hole accretion disk theory.
In section 3, we delve into the metric assumptions and Conservation Equations. We leverage the work of [13], who constructed a standard relativistic thin disk around a distorted Schwarzschild black hole up to the quadrupole. We then proceed to solve the conservation equations in the DKB metric. As the mass of the black hole is comparable to that of the sun (), the effect of self-gravity is negligible.
In section 4, we initially focus on the variations in the quadrupole moment while keeping the viscosity coefficient constant for both rotating and non-rotating configurations. Following this, we investigate a range of viscosity coefficient values for both distorted and undistorted states, ensuring that the rotational parameter remained fixed. Next, we assign specific viscosity coefficient values to the undistorted and distorted Kerr black holes, making necessary adjustments to the rotational parameter. Through detailed graphical representations, we perform a comprehensive analysis of their behaviors and the implications of these modifications. In section 5, we explain how our findings contribute to a deeper understanding of the dynamics involved in black hole configurations and their physical characteristics.
The results of our analysis, focusing on steady-state properties,such as flux,temperature, and accretion efficiency, which are derived from relativistic hydrodynamics and radiative transfer than classical thermodynamics. They are presented in the form of graphs and detailed explanations. We choose to study the DKB due to the significance of rotation in black holes and the accuracy of the model in representing the motion of the accretion disk and the central black hole. additionally, we hope that our findings will align with future observations.
Further details and explanations of the calculations are provided in Appendix A.
2 Distorted Kerr Black hole
Distortions are found in binary systems where the black hole forms an accretion disk with its compantion. They can be investigated numerically and analytically. many works have been done to numerically investigate such systems and merge black holes [20]. Also, some black holes can be modeled using axisymmetric and stationary solutions. In these cases, precise solutions for distorted black holes are examined. The Kerr metric describes the geometry of spacetime around a charge-free rotating axisymmetric black hole, which is an exact solution of Einstein’s field equations of general relativity. The DKB metric is given in the following form [11], [21]:
@@ 2 Distorted Kerr Black hole
We denote by the Legendre polynomials, which depends on the argument in all the expressions. The solution is represented in the prolate spheroidal coordinates, , and , are real constants. So:
@@
and for the spin parameter, we have:
3 Thin Disk
Accretion disks are fundamental to a wide range of astrophysical phenomena, acting as critical zones for energy release and jet formation. While significant progress has been made in understanding accretion disks around standard spherical or rotating black holes, emerging research suggests that many black holes in the universe deviate from these idealized models. Distortions caused by factors such as angular momentum distribution, gravitational wave interactions, or other astrophysical influences can profoundly impact the dynamics of surrounding accretion disks. Given this complexity, it becomes essential to establish a set of assumptions that will guide our calculations and conditions related to the DKB and its associated accretion disk:
- The disk is geometrically thin in which H is half thickness of the disk so a standard relativistic thin disk around a Kerr black hole is assumed [22] .
- It is cold which means
- The disk is optically thick which means that the mean free path of photons is short.
- We focused on the region beyond the horizon that exhibits both Axial and mass symmetry, It is rotating and this distortion maybe related to the outer parts of the accretion disk.
@@ 3.1 Conservation Equations
The equations we derive:
The first equation describes the behavior of fluid flow in a radial direction under relativistic conditions:
@@
The second equation is particles conservation:
@@
The third one is the energy conservation equation which is given as:
@@
4 Results
We work on the metric equations of curved space time and simplified them, ultimately discovering new algebraic equations as a result. The obtained results are depicted here, where we compare the plots of Kerr’s distorted and undistorted spacetime. Since the distorted solutions are only applicable in a limited region near the horizon, we focus solely on the inner portion of the ISCO disc within this spacetime. we use for slowly rotating Black hole of [25]. In the Kerr metric, the location of the event horizon depends on the spin parameter . In Boyer-Lindquist coordinates, the radius of the outer event horizon is given by [26]. For a Schwarzschild black hole (), the horizon is at , but in a distorted Kerr black hole, the horizon location varies with the spin [27]. The effectiveness of our exact solution relies heavily on the specific quadrupole selected for implementation. Shoom [10] elucidates the crucial role that this choice plays in determining the accuracy of the resulting solution. The constants involved are as follows: (All constants are in the CGS system)
@@ 4 Results
where . In the following sections, based on equations 53 to 58, we present graphs illustrating the flux, surface density, half-thickness, temperature, pressure, and viscosity as functions of the radial coordinates (distance from the black hole). To determine the event horizon of the DKB, we assumed . These graphs account for various states specified in the captions of the figures:
@@
In figures 3 and 4, the radial coordinate x is plotted against the surface density . At x = 3, we observe the maximum surface density without rotation at the initial point. However, when rotation is applied, this initial maximum shifts closer to the horizon. When we compare each quadrupole moment q in both states, we notice that in the compare rotation state (Kerr), the growth of surface density occurs more rapidly. This indicates the surface density decreases with the distance from the horizon, which we expect, but by adding rotation It is increasing that needs to be investigated more so that in higher positions it should tend to zero in order to get the correct result.
@@ 5 Discussion and conclusion
In this article, we first examined the changes in the quadrupole moment at a fixed for both rotating and non-rotating states. Next, we varied alpha for quadrupoles with values of zero, positive, and negative, while keeping the rotation parameter at a positive value less than zero. Subsequently, we fixed alpha for the zero, positive, and negative quadrupoles and adjusted the rotation parameter. By plotting all relevant graphs, we analyzed their behaviors and effects.
@@
b) -x graph: An increase in rotation (a) and a decrease in viscosity () result in an additional peak in the surface density-radial coordinate (-x) graph. A positive quadrupole accelerates the growth of the graph, while a negative quadrupole exhibits slower growth.
@@
d) T-x graph: Increasing rotation (a) and viscosity () elevate the average temperature in the temperature-radial coordinate (T-x) graph. Changes in quadrupole primarily affect the slope of the graph over time, with no noticeable impact at the initial stages.